OpenStudy (anonymous):

hellp ? Simplify quantity 1 over 9 minus 1 over x all over 1 over 81 minus 1 over x squared

OpenStudy (anonymous):

could you draw it?

OpenStudy (anonymous):

\[\frac{\frac{1}{9}-\frac{1}{x}}{\frac{1}{81}-\frac{1}{x^2}}\]

OpenStudy (anonymous):

there is a snappy way to do this, but it probably doesn't help much in understanding. multiply top and bottom of whole thing by \[81x^2\] to clear the fractions. you will get \[\frac{9x^2-81x}{x^2-81}\] as a first step. then factor numerator and deliminator

OpenStudy (anonymous):

\[\frac{9x(x-9)}{(x+9)(x-9)}\] and finally \[\frac{9x}{x+9}\] if my algebra is correct